# biblio.bib

### v1

@@ 3.4 Extension to the Brownian setting
$\mathrm{d}\overline{X}^{h,M,i}_{t}=b_{t_{h}}(\overline{X}^{h,M,i},\L(\overline{X}^{h,M,i}))\mathrm{d}t+\sigma_{t_{h}}(\overline{X}^{h,M,i},\L(\overline{X}^{h,M,i}))\mathrm{d}B^{i}_{t\wedge\tau^{i}_{M}},\quad\overline{X}^{h,M,i}_{0}=0,$
@@
$\delta^{M,i}_{t}=\int_{0}^{t}[b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))]\mathrm{d}s\\ +\int_{0}^{t}[\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))]\mathrm{d}B^{i}_{s\wedge\tau^{i}_{M}}.$
@@
$\mathbb{E}\big{[}\sup_{0\leq s\leq t}|\delta^{M,i}_{t}|^{2}\big{]}\leq 2t\int_{0}^{t}\mathbb{E}|b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))|^{2}\mathrm{d}s\\ +2\mathbb{E}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}[\sigma_{r_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{r_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))]\mathrm{d}B^{i}_{r\wedge\tau^{i}_{M}}\bigg{|}^{2}.$ (18)
@@
$b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))=b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))\\ +[b_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))].$
@@
$\mathbb{E}\big{|}b_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))\big{|}^{2}\leq C_{h}N^{-1}.$
@@
$\mathbb{E}\big{|}b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))\big{|}^{2}\leq C_{h}\mathbb{E}\big{[}\sup_{0\leq r\leq s_{h}}|X^{M,i}_{r}-\overline{X}^{M,i}_{r}|^{2}\big{]}+C_{h}N^{-1}.$
@@
$\mathbb{E}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}$ $[\sigma_{r_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{r_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))]\mathrm{d}B^{i}_{r\wedge\tau^{i}_{M}}\bigg{|}^{2}$
$\leq C_{h}\mathbb{E}\int_{0}^{t\wedge\tau^{i}_{M}}|\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))|^{2}\mathrm{d}s$
$\leq C_{h}\int_{0}^{t}\mathbb{E}|\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))|^{2}\mathrm{d}s.$
@@
$b_{s}^{M,i}:=b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))-[b_{s_{h}}(X^{\infty,i},\pi(\vec{X}^{\infty}))-b_{s_{h}}(\overline{X}^{\infty,i},\L(\overline{X}^{\infty,i}))],$
$\sigma_{s}^{M,i}:=\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))-[\sigma_{s_{h}}(X^{\infty,i},\pi(\vec{X}^{\infty}))-\sigma_{s_{h}}(\overline{X}^{\infty,i},\L(\overline{X}^{\infty,i}))],$
$\overline{\sigma}^{M,i}_{s}:=\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i})).$
@@
+\,+ $b_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))-[b_{s_{h}}(\overline{X}^{\infty,i},\pi(\vec{\overline{X}}^{\infty}))-b_{s_{h}}(\overline{X}^{\infty,i},\L(\overline{X}^{\infty,i}))].$
@@
$\mathbb{E}\big{[}\sup_{t\leq s\leq t^{\prime}}|\overline{\sigma}^{M,i}_{s}|^{2}\big{]}\leq 2\mathbb{E}\big{[}\sup_{0\leq s\leq T}|\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))|^{2}\big{]}+\\ 2\mathbb{E}\big{[}\sup_{0\leq s\leq T}|\sigma_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\L(\overline{X}^{M,i}))|^{2}\big{]}.$
@@
$\mathrm{d}\delta X^{M}_{t_{h}}$ $=\big{[}D_{x}\sigma_{t_{h}}(X^{M},\L(X^{M}))\cdot\delta X^{M}+\delta_{P}\sigma_{t_{h}}(X^{M},\L(X^{M}),\tilde{X}^{h})\big{]}\mathrm{d}B_{t\wedge\tau_{M}}$
$+\big{[}D_{x}b_{t_{h}}(X^{M},\L(X^{M}))\cdot\delta X^{M}+\delta_{P}b_{t_{h}}(X^{M},\L(X^{M}),\tilde{X}^{M})\big{]}\mathrm{d}t$
$+\mathbb{E}\big{[}D_{y}\delta_{P}b_{t_{h}}(X^{M},\L(X^{M}),X^{M})\cdot\delta X^{M}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}t$
$+\mathbb{E}\big{[}D_{y}\delta_{P}\sigma_{t_{h}}(X^{M},\L(X^{M}),X^{M})\cdot\delta X^{M}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}B_{t\wedge\tau_{M}},\qquad\delta X^{M}_{0}=0.\phantom{abcd}$
@@ 4.1 Large deviations and stochastic control
$J^{h}_{\Sigma}(\vec{X}^{h,N}_{0},\vec{u}^{N}):=N^{-1}H(\L(\vec{X}^{h,N}_{0})|\L(X^{1}_{0})^{\otimes N})+\mathbb{E}\big{[}\mathcal{E}(\vec{u}^{N})+F(\pi(\vec{X}^{h,N,\vec{u}^{N}}))\big{]},$
@@
    $\mathbb{E}[|X^{h,1,N}_{0}|^{p}]\leq C$, where $C$ only depends on $(p,\lVert F\rVert_{\infty},\L(X^{1}_{0}))$.
@@
By choosing Ł$\L(\vec{X}^{h,N}_{0})=\L(X^{1}_{0})^{\otimes N}$ and $\vec{u}^{N}\equiv 0$, we get that $\mathcal{F}^{h,N}\leq\|F\|_{\infty}.$ This allows us to restrict ourselves to admissible controls $\vec{u}^{N}$ with $\mathbb{E}[\mathcal{E}(\vec{u}^{N})]\leq 2\|F\|_{\infty}$. Let us fix such a $(\vec{X}^{h,N}_{0},\vec{u}^{N})$. Let $\tau$ be a uniformly distributed random permutation of $\{1,\ldots,N\}$ that is independent of $\vec{B}^{N}$, $\vec{X}_{0}^{h,N}$, and $\vec{u}^{N}$. Defining $X^{h,\tau,i,N}_{0}:=X^{h,\tau(i),N}_{0}$ and $u^{\tau,i,N}:=u^{\tau(i),N}$, we notice that $(X^{h,\tau,i,N}_{0},u^{\tau,i,N})_{1\leq i\leq N}$ is exchangeable. The symmetry of (22) then implies that $\pi(X^{h,\tau,N,\vec{u}^{\tau,N}})$ has the same law as $\pi(X^{h,N,\vec{u}^{N}})$. For any permutation $\tau^{\prime}$, we define
@@
$H(\L(\vec{X}^{h,\tau^{\prime},N}_{0})|\L(X^{1}_{0})^{\otimes N})=H((f^{N}_{\tau^{\prime}})_{\#}\L(\vec{X}^{h,N}_{0})|(f^{N}_{\tau^{\prime}})_{\#}\L(X^{1}_{0})^{\otimes N})\leq H(\L(\vec{X}^{h,N}_{0})|\L(X^{1}_{0})^{\otimes N}).$
Since Ł$\L(\vec{X}^{h,\tau,N}_{0})=\mathbb{E}[(f^{N}_{\tau})_{\#}\L(\vec{X}^{h,N}_{0})]$ and $P\mapsto H(P|Q)$ is convex, Jensen’s inequality then implies that
$H(\L(\vec{X}^{h,\tau,N}_{0})|\L(X^{1}_{0})^{\otimes N})\leq\mathbb{E}[H((f^{N}_{\tau})_{\#}\L(\vec{X}^{h,N}_{0})|\L(X^{1}_{0})^{\otimes N})]\leq H(\L(\vec{X}^{h,N}_{0})|\L(X^{1}_{0})^{\otimes N}).$
@@
$N^{-1}H(\L(\vec{X}^{h,N}_{0})|\L(X^{1}_{0})^{\otimes N})\leq 2\lVert F\rVert_{\infty}.$
@@
$H(\L(X^{h,1,N}_{0})|\L(X^{1}_{0}))\leq N^{-1}H(\L(\vec{X}^{h,N}_{0})|\L(X^{1}_{0})^{\otimes N}).$
@@
$H(\L(X^{h,1,N}_{0})|\L(X^{1}_{0}))=\sup_{\begin{subarray}{c}\phi\text{ measurable}\\ \mathbb{E}[\phi(X^{1}_{0})]<+\infty\end{subarray}}\mathbb{E}[\phi(X^{h,1,N}_{0})]-\log\mathbb{E}[e^{\phi(X^{1}_{0})}],$ (23)
@@
$\mathcal{F}^{h,N}=\inf_{R_{N}\in\mathcal{P}_{p}((\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})^{N})}N^{-1}H(R_{N}|\L(X^{1}_{0},B^{1})^{\otimes N})+N^{-1}\int_{\mathbb{R}^{d}\times\mathcal{C}^{d}}[F\circ\Psi_{h}]\mathrm{d}R_{N}.$
@@
For $h\in(0,1]$, Proposition 2.7 gives that the sequence of the Ł$\L(\Psi_{h}(\pi(\vec{X}^{N}_{0},\vec{B}^{N})))$ satisfies the LDP with good rate function $P\in\mathcal{P}_{p}(\mathcal{C}^{d})\mapsto H(P|\Gamma_{h}(P))$. Varadhan’s lemma [dupuis2011weak, Theorem 1.2.1] then gives that
@@
for every bounded Lipschitz $F:\mathcal{P}_{p}(\mathcal{C}^{d})\rightarrow\mathbb{R}$. From [dupuis2011weak, Corollary 1.2.5], this proves that the sequence of the Ł$\L(\pi(\vec{X}^{0,N}))$ satisfies the LDP with good rate function $P\mapsto H(P|\Gamma_{0}(P))$, this rate function being indeed good from Corollary B.4. The $\Gamma$-convergence of $I_{h}$ towards $I_{0}$ precisely corresponds to the proof of Proposition B.6 with $F\equiv 0$. ∎

### v2

@@ 3.4 Extension to the Brownian setting
$\mathrm{d}\overline{X}^{h,M,i}_{t}=b_{t_{h}}(\overline{X}^{h,M,i},\mathcal{L}(\overline{X}^{h,M,i}))\mathrm{d}t+\sigma_{t_{h}}(\overline{X}^{h,M,i},\mathcal{L}(\overline{X}^{h,M,i}))\mathrm{d}B^{i}_{t\wedge\tau^{i}_{M}},\quad\overline{X}^{h,M,i}_{0}=0,$
@@
$\delta^{M,i}_{t}=\int_{0}^{t}[b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))]\mathrm{d}s\\ +\int_{0}^{t}[\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))]\mathrm{d}B^{i}_{s\wedge\tau^{i}_{M}}.$
@@
$\mathbb{E}\big{[}\sup_{0\leq s\leq t}|\delta^{M,i}_{t}|^{2}\big{]}\leq 2t\int_{0}^{t}\mathbb{E}|b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))|^{2}\mathrm{d}s\\ +2\mathbb{E}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}[\sigma_{r_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{r_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))]\mathrm{d}B^{i}_{r\wedge\tau^{i}_{M}}\bigg{|}^{2}.$ (18)
@@
$b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))=b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))\\ +[b_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))].$
@@
$\mathbb{E}\big{|}b_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))\big{|}^{2}\leq C_{h}N^{-1}.$
@@
$\mathbb{E}\big{|}b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))\big{|}^{2}\leq C_{h}\mathbb{E}\big{[}\sup_{0\leq r\leq s_{h}}|X^{M,i}_{r}-\overline{X}^{M,i}_{r}|^{2}\big{]}+C_{h}N^{-1}.$
@@
$\mathbb{E}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}$ $[\sigma_{r_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{r_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))]\mathrm{d}B^{i}_{r\wedge\tau^{i}_{M}}\bigg{|}^{2}$
$\leq C_{h}\mathbb{E}\int_{0}^{t\wedge\tau^{i}_{M}}|\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))|^{2}\mathrm{d}s$
$\leq C_{h}\int_{0}^{t}\mathbb{E}|\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))|^{2}\mathrm{d}s.$
@@
$b_{s}^{M,i}:=b_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))-[b_{s_{h}}(X^{\infty,i},\pi(\vec{X}^{\infty}))-b_{s_{h}}(\overline{X}^{\infty,i},\mathcal{L}(\overline{X}^{\infty,i}))],$
$\sigma_{s}^{M,i}:=\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))-[\sigma_{s_{h}}(X^{\infty,i},\pi(\vec{X}^{\infty}))-\sigma_{s_{h}}(\overline{X}^{\infty,i},\mathcal{L}(\overline{X}^{\infty,i}))],$
$\overline{\sigma}^{M,i}_{s}:=\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i})).$
@@
+\,+ $b_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))-b_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))-[b_{s_{h}}(\overline{X}^{\infty,i},\pi(\vec{\overline{X}}^{\infty}))-b_{s_{h}}(\overline{X}^{\infty,i},\mathcal{L}(\overline{X}^{\infty,i}))].$
@@
$\mathbb{E}\big{[}\sup_{t\leq s\leq t^{\prime}}|\overline{\sigma}^{M,i}_{s}|^{2}\big{]}\leq 2\mathbb{E}\big{[}\sup_{0\leq s\leq T}|\sigma_{s_{h}}(X^{M,i},\pi(\vec{X}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))|^{2}\big{]}+\\ 2\mathbb{E}\big{[}\sup_{0\leq s\leq T}|\sigma_{s_{h}}(\overline{X}^{M,i},\pi(\vec{\overline{X}}^{M}))-\sigma_{s_{h}}(\overline{X}^{M,i},\mathcal{L}(\overline{X}^{M,i}))|^{2}\big{]}.$
@@
$\mathrm{d}\delta X^{M}_{t_{h}}$ $=\big{[}D_{x}\sigma_{t_{h}}(X^{M},\mathcal{L}(X^{M}))\cdot\delta X^{M}+\delta_{P}\sigma_{t_{h}}(X^{M},\mathcal{L}(X^{M}),\tilde{X}^{h})\big{]}\mathrm{d}B_{t\wedge\tau_{M}}$
$+\big{[}D_{x}b_{t_{h}}(X^{M},\mathcal{L}(X^{M}))\cdot\delta X^{M}+\delta_{P}b_{t_{h}}(X^{M},\mathcal{L}(X^{M}),\tilde{X}^{M})\big{]}\mathrm{d}t$
$+\mathbb{E}\big{[}D_{y}\delta_{P}b_{t_{h}}(X^{M},\mathcal{L}(X^{M}),X^{M})\cdot\delta X^{M}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}t$
$+\mathbb{E}\big{[}D_{y}\delta_{P}\sigma_{t_{h}}(X^{M},\mathcal{L}(X^{M}),X^{M})\cdot\delta X^{M}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}B_{t\wedge\tau_{M}},\qquad\delta X^{M}_{0}=0.\phantom{abcd}$
@@ 4.1 Large deviations and stochastic control
$J^{h}_{\Sigma}(\vec{X}^{h,N}_{0},\vec{u}^{N}):=N^{-1}H(\mathcal{L}(\vec{X}^{h,N}_{0})|\mathcal{L}(X^{1}_{0})^{\otimes N})+\mathbb{E}\big{[}\mathcal{E}(\vec{u}^{N})+F(\pi(\vec{X}^{h,N,\vec{u}^{N}}))\big{]},$
@@
    $\mathbb{E}[|X^{h,1,N}_{0}|^{p}]\leq C$, where $C$ only depends on $(p,\lVert F\rVert_{\infty},\mathcal{L}(X^{1}_{0}))$.
@@
By choosing $\mathcal{L}(\vec{X}^{h,N}_{0})=\mathcal{L}(X^{1}_{0})^{\otimes N}$ and $\vec{u}^{N}\equiv 0$, we get that $\mathcal{F}^{h,N}\leq\|F\|_{\infty}.$ This allows us to restrict ourselves to admissible controls $\vec{u}^{N}$ with $\mathbb{E}[\mathcal{E}(\vec{u}^{N})]\leq 2\|F\|_{\infty}$. Let us fix such a $(\vec{X}^{h,N}_{0},\vec{u}^{N})$. Let $\tau$ be a uniformly distributed random permutation of $\{1,\ldots,N\}$ that is independent of $\vec{B}^{N}$, $\vec{X}_{0}^{h,N}$, and $\vec{u}^{N}$. Defining $X^{h,\tau,i,N}_{0}:=X^{h,\tau(i),N}_{0}$ and $u^{\tau,i,N}:=u^{\tau(i),N}$, we notice that $(X^{h,\tau,i,N}_{0},u^{\tau,i,N})_{1\leq i\leq N}$ is exchangeable. The symmetry of (22) then implies that $\pi(X^{h,\tau,N,\vec{u}^{\tau,N}})$ has the same law as $\pi(X^{h,N,\vec{u}^{N}})$. For any permutation $\tau^{\prime}$, we define
@@
$H(\mathcal{L}(\vec{X}^{h,\tau^{\prime},N}_{0})|\mathcal{L}(X^{1}_{0})^{\otimes N})=H((f^{N}_{\tau^{\prime}})_{\#}\mathcal{L}(\vec{X}^{h,N}_{0})|(f^{N}_{\tau^{\prime}})_{\#}\mathcal{L}(X^{1}_{0})^{\otimes N})\leq H(\mathcal{L}(\vec{X}^{h,N}_{0})|\mathcal{L}(X^{1}_{0})^{\otimes N}).$
Since $\mathcal{L}(\vec{X}^{h,\tau,N}_{0})=\mathbb{E}[(f^{N}_{\tau})_{\#}\mathcal{L}(\vec{X}^{h,N}_{0})]$ and $P\mapsto H(P|Q)$ is convex, Jensen’s inequality then implies that
$H(\mathcal{L}(\vec{X}^{h,\tau,N}_{0})|\mathcal{L}(X^{1}_{0})^{\otimes N})\leq\mathbb{E}[H((f^{N}_{\tau})_{\#}\mathcal{L}(\vec{X}^{h,N}_{0})|\mathcal{L}(X^{1}_{0})^{\otimes N})]\leq H(\mathcal{L}(\vec{X}^{h,N}_{0})|\mathcal{L}(X^{1}_{0})^{\otimes N}).$
@@
$N^{-1}H(\mathcal{L}(\vec{X}^{h,N}_{0})|\mathcal{L}(X^{1}_{0})^{\otimes N})\leq 2\lVert F\rVert_{\infty}.$
@@
$H(\mathcal{L}(X^{h,1,N}_{0})|\mathcal{L}(X^{1}_{0}))\leq N^{-1}H(\mathcal{L}(\vec{X}^{h,N}_{0})|\mathcal{L}(X^{1}_{0})^{\otimes N}).$
@@
$H(\mathcal{L}(X^{h,1,N}_{0})|\mathcal{L}(X^{1}_{0}))=\sup_{\begin{subarray}{c}\phi\text{ measurable}\\ \mathbb{E}[\phi(X^{1}_{0})]<+\infty\end{subarray}}\mathbb{E}[\phi(X^{h,1,N}_{0})]-\log\mathbb{E}[e^{\phi(X^{1}_{0})}],$ (23)
@@
$\mathcal{F}^{h,N}=\inf_{R_{N}\in\mathcal{P}_{p}((\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})^{N})}N^{-1}H(R_{N}|\mathcal{L}(X^{1}_{0},B^{1})^{\otimes N})+N^{-1}\int_{\mathbb{R}^{d}\times\mathcal{C}^{d}}[F\circ\Psi_{h}]\mathrm{d}R_{N}.$
@@
For $h\in(0,1]$, Proposition 2.7 gives that the sequence of the $\mathcal{L}(\Psi_{h}(\pi(\vec{X}^{N}_{0},\vec{B}^{N})))$ satisfies the LDP with good rate function $P\in\mathcal{P}_{p}(\mathcal{C}^{d})\mapsto H(P|\Gamma_{h}(P))$. Varadhan’s lemma [dupuis2011weak, Theorem 1.2.1] then gives that
@@
for every bounded Lipschitz $F:\mathcal{P}_{p}(\mathcal{C}^{d})\rightarrow\mathbb{R}$. From [dupuis2011weak, Corollary 1.2.5], this proves that the sequence of the $\mathcal{L}(\pi(\vec{X}^{0,N}))$ satisfies the LDP with good rate function $P\mapsto H(P|\Gamma_{0}(P))$, this rate function being indeed good from Corollary B.4. The $\Gamma$-convergence of $I_{h}$ towards $I_{0}$ precisely corresponds to the proof of Proposition B.6 with $F\equiv 0$. ∎